Grade 11IIT JEE Entrance Exam

Explain the hybridization in N2 ......!..(.............................

Sudhanva G V
7 Years agoGrade 11

1 Answer

Arun
7 Years ago

for nitrogen , considering any one nitrogen atom . Each nitrogen atom is attached to one more nitrogen atom , each nitrogen atom contains 5 valence electrons and is forming a triple covalent bond , so the valency of each nitrogen atom is 3 . Also the molecule contains no charge that is the charge over the molecule is zero . So putting the values we get ,

H = 1 + (5 - 3 - 0)/2

H = 1 + 2/2

H = 1 + 1

H= 2

So each atom in nitrogen molecule is having hybridization number equal to 2 or a hybridization state of sp .